### v1

# SplitVAEs: Decentralized scenario generation from siloed data for stochastic optimization problems

Abstract
Stochastic optimization problems in large-scale multi-stakeholder networked systems (e.g., power grids and supply chains) rely on data-driven scenarios to encapsulate complex spatiotemporal interdependencies. However, centralized aggregation of stakeholder data is challenging due to the existence of data silos resulting from computational and logistical bottlenecks. In this paper, we present SplitVAEs, a decentralized scenario generation framework that leverages variational autoencoders to generate high-quality scenarios without moving stakeholder data. With the help of experiments on distributed memory systems, we demonstrate the broad applicability of SplitVAEs in a variety of domain areas that are dominated by a large number of stakeholders. Our experiments indicate that SplitVAEs can learn spatial and temporal interdependencies in large-scale networks to generate scenarios that match the joint historical distribution of stakeholder data in a decentralized manner. Our experiments show that SplitVAEs deliver robust performance compared to centralized, state-of-the-art benchmark methods while significantly reducing data transmission costs, leading to a scalable, privacy-enhancing alternative to scenario generation.
Index Terms:
Scenario generation, stochastic optimization, decentralized learning, variational autoencoders.
I Introduction
Stochastic optimization (SO) problems are essential components in large-scale, data-driven planning and decision frameworks for multi-stakeholder-led networked systems such as energy [1, 2], supply chain [3], spare parts management [4], healthcare [5] and logistics [6]. Naturally, good quality solutions to SO problems require high-fidelity data-driven scenarios that encapsulate spatially and temporally correlated interdependencies from myriad sources. As a result, generating high-fidelity scenarios depends on the availability of high-quality datasets that capture these interdependencies from heterogeneous data sources. However, conventional scenario generation methods necessitate centralized data aggregation to capture interdependencies to drive stochastic planning decisions. Centralized aggregation schemes for scenario generation face several computational, and logistical challenges [7, 8] due to the presence of data silos across multiple stakeholders of a networked system. Therefore, in this study, we present SplitVAEs, a novel decentralized scenario generation framework that utilizes variational autoencoders (VAE) to eliminate centralized data aggregation. SplitVAEs are compatible with pre-existing data silos, ensure scalable generation of high-fidelity scenarios and significantly reduce data transfer costs.
Conventionally the process of solving an SO problem begins with the acquisition of local spatiotemporally correlated datasets from individual stakeholders. The mechanism of data aggregation is typically carried out by a control center or planning organization as represented in Figure 1. Such coordinating entities are typically found in multi-stakeholder networked systems such as Independent System Operator (ISO) in power systems and control towers in the supply chain. The aggregated dataset is then used to generate high-fidelity scenarios that can be used to solve an SO problem in a centralized fashion at the control center level. The solutions to these SO problems represent data-driven planning decisions that can be in turn used to determine an operational schedule for each stakeholder subject to constraints like reliability, safety and demand satisfaction. However, the entire data-driven planning mechanism is challenged in cases where data silos are present across various stakeholders.
Figure 1: Scenario generation for stochastic optimization.
Therefore, the goal of SplitVAEs is to enable planning organization and control centers to generate generate stochastic decisions without incurring the burden of data processing and aggregation from multiple system stakeholders. Such planning organizations typically have regulatory oversight on the operational aspects of these systems necessitating SO problems to be solved in a high-frequency fashion. Thus, our proposed SplitVAEs framework relies on distributed memory-based aggregation primitives to conduct training without the need to move local data. As a result, the SplitVAEs algorithmic framework, by its very design is fully capable of functioning in the presence of geographically distributed, decentralized data silos as is typically found in multi-stakeholder infrastructure systems. Therefore, the SplitVAEs framework can potentially be extended to operate in hybrid edge-cloud environments based on containerized applications enabling significant reductions in data transmission. Additionally, it is capable of succinctly capturing spatiotemporal interdependencies in a decentralized manner despite the inherent non-linearities in the underlying data among multiple stakeholder subsystems while maintaining high scalability. SplitVAEs complement existing stochastic optimization frameworks by supplying high-fidelity scenarios generated from siloed datasets. More importantly, SplitVAEs are not meant to be a replacement of optimization solution methodologies. As a result, we can concisely summarize the contributions of our study as follows:
  • We design a computationally efficient learning paradigm that combines edge-based autoencoders with a server-driven variational autoencoder to jointly learn spatiotemporal interdependencies of siloed data.
  • We develop a scalable algorithmic framework that decomposes global backpropagation steps into edge and server-based sub-problems, enabling the bi-directional flow of learning insights, eliminating the movement of data.
  • We propose a distributed memory-based computational paradigm that can be seamlessly adopted to enable a scalable, real-world implementation of the SplitVAEs framework to generate scenarios for SO problems.
  • We demonstrate the ability of SplitVAEs to handle heterogeneous datasets across diverse areas by comparing the generated scenarios with established benchmark methods.
The remaining parts of this study are organized in the following manner. Section II provides a literature review of prior work. Sections III and IV describe the architecture and the algorithmic design of our proposed framework, respectively. Section V provides a summary of benchmark methods as well as statistical evaluation metrics. Section VI analyze the performance of the SplitVAEs framework across different application domains and diverse computational settings. Section VII serves as the concluding section of this study.
II Related Works
There have been several scenario generation mechanisms that have been proposed including auto-regressive moving average (ARIMA) methods [9, 1], Gaussian copula-based approaches [10, 11] and deep generative adversarial networks (GANs) [12, 7]. A majority of these studies, however, are centralized in nature leading to computational and data privacy related concerns [7]. From a computational standpoint, centralized methods such as ARIMA and copula-based approaches rely on the construction and factorization of covariance matrices [13, 11] aided by the accumulation of stakeholder data. On the other hand, centralized GAN-based methods can be notoriously difficult to train owing to high instability and challenges in convergence [14]. As a result, scalability remains an issue with these methods, especially for large-scale networks that consist of many sources of uncertainty.
Centralized methods also depend on nonlinear modeling of relationships between predictor feature variables and the generated scenarios are connected in space and time. The authors in [13] used wind speed data to construct scenarios of wind power generation via the power curve of wind turbines, while the work in [15] leveraged the nonlinear relationship between temperature and power consumption using epi-spline functions to generate useful scenarios. These approaches might become untenable, especially when integrating numerous heterogeneous sources of data, including but not limited to wind, PV farms, and alternative fuel sources such as hydrogen.
On the other hand, VAE [16] serve as a potential alternative to methods that rely on non-linear modeling requirements. VAEs can help synthesize new, previously unseen scenarios by capturing the distribution of more concise, low-dimensional embeddings in a latent space. By sampling from the learned latent space distributions, VAEs can be used to generate scenarios that are spatially and temporally interdependent. Unlike GANs however, the VAEs provide a scalable, and decomposable implementation paradigm which is suited for data silos with heterogenous features as well.
Different application areas have their own challenges with respect to siloed data. For example, visibility across stakeholder subsystems is a major hindrance in supply chain and logistics owing to heterogeneity and incompatibility of local data storage systems [17]. Similarly, in the medical domain, data privacy is the primary reason for the existence of data silos [18]. Privacy obstacles to scenario generation could also stem from the sensitivity of underlying datasets from the commercial standpoint [19]. Such data sharing challenges are especially prevalent in the case of variable renewable energy (VRE) stakeholders [20, 8]. Therefore, conventional methods of scenario generation face several limitations in terms of their applicability to heterogeneous data from diverse stakeholders.
Split learning [21, 22, 23] presents a more generalized approach to handling siloed data sources in a decentralized, privacy-friendly fashion. Instead of relying on horizontal partitioning of data, split learning relies on model partitioning schemes that yield low-dimensional embeddings from heterogeneous data. Split learning paradigms leverage these embeddings to learn the interdependencies of the features contained in the siloed heterogeneous datasets without the need to move local data. As a result, split learning methods can potentially help resolve the data heterogeneity and data silo challenges inherent in scenario generation methods. Therefore, our approach combines the split learning paradigm with the VAE to yield a computationally efficient, privacy-preserving framework capable of delivering high-quality scenarios from decentralized and siloed datasets.
Figure 2: The learning paradigm of SplitVAEs.
III SplitVAEs Framework Design
We consider the SplitVAEs framework from the perspective of edge and server-level agents. The edge can be used to represent heterogeneous system stakeholders that govern individual data silos comprising time-series datasets. In our framework, the server houses the VAE and represents the control center or planning organization, while the edge entities consist of an autoencoder (AE) that is jointly trained in conjunction with the VAE at the server level. We begin our discussion by describing the conventional, centralized VAE (Central-VAE) formulation.
III-A Centralized variational autoencoder (VAE) formulation
The Central-VAE model consumes raw data $y\in\mathbb{R}^{d}$ to yield latent space embeddings, denoted by $z\in\mathbb{R}^{s}$, with the help of a probabilistic encoder $q^{cvae}_{\phi}(z|y)$. Specifically, $q^{cvae}_{\phi}(z|y)$ initially predicts the mean and variance $(\mu_{s},\sigma_{s})$ of the latent space distribution as represented in Equation (1). The $(\mu_{s},\sigma_{s})$ are used to generate latent space representations $z$ with the help of sampling from the standard normal distribution as represented in Equation (2). The integration of a random sample $\epsilon$ is referred to as the reparametrization trick and is necessary to avoid computationally expensive simulation-driven latent space estimation. Similarly, a probabilistic decoder $p^{cvae}_{\theta}(y|z)$ is applied so as to obtain a reconstruction $\tilde{y}\in\mathbb{R}^{d}$ of the raw data $y$ represented by Equation (3).
$(\mu,\sigma)\leftarrow q^{cvae}_{\phi}(x)$ (1)
$z\leftarrow\mu+\sigma\odot\epsilon_{i},\text{ where }\epsilon_{i}\sim\mathcal{N}(0,I)$ (2)
$\tilde{x}_{i}\leftarrow p^{cvae}_{\theta}(z_{i})$ (3)
Equations (1) to (3) are considered to be one forward pass of the VAE model. On the other hand, the backward pass of the VAE is carried out with the help of a total loss function represented by Equation (4). The total loss comprises of two errors: the reconstruction error, denoted as $L^{recon}(\theta,x,z)$ and indicated by Equation (5); the Kullback-Leibler divergence, denoted as $L^{KL}(\phi,\theta,x,z)$ and derived in (6), which measures the similarity of latent space distributions based on the probabilistic encoders and decoders.
$L(\theta,\phi,x,z)=L^{recon}(\theta,x,z)+L^{KL}(\phi,\theta,x,z)\text{\quad}$ (4)
$L^{recon}(\theta,x,z)=-\mathbb{E}_{z\sim q^{cvae}_{\phi}(z|x)}\Big{(}\log p^{cvae}_{\theta}(x|z)\Big{)}\text{ }$ (5)
$L^{KL}(\phi,\theta,x,z)=D_{KL}\Big{(}q^{cvae}_{\phi}(z|x)||p^{cvae}_{\theta}(z)\Big{)}\text{\quad\quad}$ (6)
Therefore, the implementation of the Central-VAE requires the transfer of edge-level stakeholder datasets to compute the reconstruction loss as well as generate latent space embeddings, which might be infeasible due to the reasons outlined in Section I. Therefore, to learn the spatiotemporal interdependencies despite the decentralized, siloed nature of stakeholder data, the SplitVAEs framework incorporates a novel split training mechanism that leverages an edge-based AE component and a server-based VAE component.
III-B Edge-level autoencoder (AE) formulation
Given the aforementioned reasons, at the edge-level, we implement an autoencoder (AE) framework comprising deep neural network (DNN) encoder and decoder models. We denote the siloed time-series data located at each edge location $n\in\mathcal{N}$ as $y^{n}$. The functional forms of the encoder and decoder components are denoted by $x^{n}\leftarrow F_{\theta^{n}_{E}}(y^{n})$ and $\tilde{y}^{n}\leftarrow G_{\theta^{n}_{D}}(x^{n})$, respectively. For the edge-based AE framework, let $x^{n},\tilde{y}^{n},\theta^{n}_{E},\theta^{n}_{D}$ be the low-dimensional embedding, the reconstructed vector, and the model parameters of the encoder and decoder, respectively. Unlike conventional AE training, a unique aspect of the SplitVAEs training mechanism is the independent training of the encoder and decoder components to holistically capture spatial and temporal interdependencies among all the edge devices.
III-C Server-level VAE formulation
At the server level, we implement a conventional VAE framework. However, unlike the Central-VAE, in each training epoch, the server model consumes the set of low dimensional embeddings, denoted as $\{x^{\{1\}},\ldots,x^{\{N\}}\}$, supplied by the edge-based AEs at the various stakeholders. Let $q_{\phi}(z|x)$ be the probabilistic encoder, responsible for deriving the latent vector $z$ from the input embeddings $x$ in terms of empirical mean $\mu$ and standard deviation $\sigma$, respectively. Additionally, we denote $z\leftarrow R_{\mu,\sigma}(\epsilon)$ as the reparametrization function akin to Equation (2). The second core component is the latent vector $z$ representing the compressed data in a lower dimension. The size of $z$ becomes a crucial hyperparameter that necessitates careful tuning, in conjunction with the overall architecture of the entire network. The final component is the probabilistic decoder, denoted as $p_{\theta}(x|z)$, responsible for reconstructing a data point $x$ from a given latent vector $z$.
III-D Loss Functions
Here, we discuss the relevant formulations of the reconstruction error and the KL loss that are utilized by the SplitVAEs framework for training.
Reconstruction Loss: We employ the binary cross-entropy (BC) loss function at the $n$-th edge location for all $n\in\mathcal{N}$ to measure the mean reconstruction error between predicted values $(\tilde{y})^{n}$ and observed values $(y)^{n}$ as represented in Equation (7).
$L^{BC}_{n}=-\frac{1}{B}\sum\limits_{b=1}^{B}\Big{[}\tilde{y}^{b}log(y^{b})+(1-\tilde{y}^{b})log(1-y^{b})\Big{]}$ (7)
In Equation (7), $B,\tilde{y}_{b},y_{b}$ represent the batch size, predicted, and target data for scenario generation, respectively. To preserve the simplicity of notations, we assume without loss of generality that $L^{BC}_{n}$ is computed concerning flattened representations of local tensor data and estimates $(y)^{n},(\tilde{y})^{n}$ of dimension $d^{n}$. We note that the computation of $L^{BC}_{n}$ can be carried out independently at each edge location since it relies purely on siloed data $(y)^{n}$.
KL Loss: We utilize the formulation given in Equation (8) to compute the KL loss as follows:
$L^{KL}=-\frac{1}{2}\Big{[}1+2\log(\sigma)-\mu^{2}-\sigma\Big{]}$ (8)
In Equation (8), the KL loss depends only on the latent space embeddings that result from the reparametrization trick described in Section III-A. Consequently, the backpropagation of the error resulting from the KL loss can be initiated from the server level itself. Therefore, the decomposed reconstruction and KL error terms can be leveraged to effect a computationally efficient training mechanism of the SplitVAEs framework.
IV SplitVAEs Algorithmic Framework
There are two crucial characteristics of the SplitVAEs decomposition scheme that can be advantageous to its training. First, the KL loss, $L^{KL}$, is purely a factor of the latent space embeddings estimated by the server-based VAE framework. Second, the reconstruction error can be computed in a fully decentralized fashion by each of the edge locations locally without revealing the siloed datasets themselves. Thus, only the local reconstruction loss needs to be back-propagated from the edge-based AE decoder to the server-based VAE decoder. This step can be immediately followed by a backpropagation of the total loss, $L^{recon}+L^{KL}$, from the server-level VAE encoder to the edge-based AE encoders. As a result of our two-pronged backpropagation mechanism, we can enable the flow of insights back and forth between edge and server-level devices, leading to seamless learning of the overall spatiotemporal interdependencies among stakeholders.
A single training epoch in our proposed framework can be divided based on edge- and server-level algorithmic frameworks. In the following sections, we discuss in detail the decomposition of the respective frameworks about the forward and backward passes of the respective model components, i.e., the server-level VAE and the encoders and decoders of AE. Specifically, we consider a set of $\mathcal{N}$ edge nodes wherein $|\mathcal{N}|=N$. To accomplish the SplitVAEs training, we leverage several common distributed memory aggregation primitives such as Scatter and Gather, denoted as DM.Scatter and DM.Gather, respectively. Tensor.Concat, Tensor.Split, BACKPROP functions indicate tensor concatenation, split and backpropagation primitives respectively. We designate the server as the root process $r$, while the edge-level processes are denoted by $n$ similar to the concept of ranks in distributed memory frameworks. We use $k$ to denote iterative updates of the model parameters represented by $\theta^{n}_{E},\theta^{n}_{D},\phi,\theta$.
IV-A Edge-based Algorithmic Framework
One training epoch of the edge-based AE framework consists of the following computational steps.
AE Encoder Forward Pass: The computational steps associated with the forward pass of the AE encoder are represented by the function EdgeEncFP. In EdgeEncFP, we consider local siloed training data $y^{b}_{e}$ which is passed through the AE encoder $F_{(\theta^{n}_{E})^{k}}$ to yield the low dimensional embedding $(x^{b}_{e})^{n}$. We utilize a DM.Gather operation to aggregate the embeddings at the server level.
Algorithm 1 AE Encoder Forward Pass
function EncEdgeFP(root:r,agent:n,epoch:$e$,batch:$b$)
  acquire local data $y^{b}_{e}$
  obtain $(x^{b}_{e})^{n}\leftarrow F_{(\theta^{n}_{E})^{k}}(y^{b}_{e})$
  DM.Gather(root:r,data:$(x^{b}_{e})^{n}$)
end function
AE Decoder Forward Pass: The function EdgeDecFP represents the reconstruction of siloed edge data. We leverage a DM.Scatter operation to collect the predicted values of the low-dimensional embedding outcomes of the server-level training denoted by $(\tilde{x}^{b}_{e})^{n}$. These predictions are passed through the AE decoder $G_{\theta^{k}_{D}}$ to generate reconstructions and the BCE loss through Equation (7).
Algorithm 2 AE Decoder Forward Pass
function EdgeDecFP(root:r,agent:n,epoch:$e$,batch:$b$)
  $(\tilde{x}^{b}_{e})^{n}\leftarrow$DM.Scatter(root:r,data:NULL)
  obtain $\tilde{y}^{b}_{e}\leftarrow G_{\theta^{k}_{D}}(\tilde{x}^{b}_{e})$
  compute reconstruction loss $({L_{BCE}})^{b}_{e}$ using (7)
end function
AE Decoder Backward Pass: As an immediate next step, our implementation leverages EdgeDecBP to back-propagate the errors based on the reconstruction losses. In EdgeDecBP $(g_{\theta_{D}})^{b}_{e}=\nabla_{\theta_{D}}(L^{recon}_{n})^{b}_{e}$ represents the gradient of $(L^{recon}_{n})^{b}_{e}$ concerning $\theta_{D}$ and $\Delta(\tilde{x})^{b}_{e}$ represents the error at the input layer of the edge-based decoder. Finally, we use a DM.Gather operation to collect the errors of the low-dimensional embeddings $(\tilde{x}^{b}_{e})^{n}$ at the server.
Algorithm 3 AE Decoder Backward Pass
function EdgeDecBP(root:r,agent:n,epoch:$e$,batch:$b$)
  obtain $(\Delta\tilde{x})^{b}_{e},(g_{D})^{b}_{e}\leftarrow$ BACKPROP($G_{(\theta^{n}_{D})^{k}},(L^{BC}_{n})^{b}_{e}$)
  update $\theta^{k+1}_{D}\leftarrow\theta^{k}_{D}-\eta_{D}.(g_{D})^{b}_{e}$
  DM.Gather(root:r,data:$(\Delta\tilde{x}^{b}_{e})^{n}$)
end function
AE Encoder Backward Pass: The last step of the edge based training epoch is presented in EdgeEncBP. We leverage DM.Scatter operation that collects errors of low dimensional embeddings $(\Delta{x}^{b}_{e})^{n}$ computed by the server. Consequently, the AE encoder back-propagates these errors through the local encoder $F_{(\theta^{n}_{E})^{k}}$ to complete the training epoch.
Algorithm 4 AE Encoder Backward Pass
function EdgeEncBP(root:r,agent:n,epoch:$e$,batch:$b$)
  $(\Delta{x}^{b}_{e})^{n}\leftarrow$DM.Scatter(root:r,data:NULL)
  $(\Delta y)^{b}_{e},(g_{E})^{b}_{e}\leftarrow$BACKPROP($F_{(\theta^{n}_{E})^{k}},\Delta{x}^{b}_{e}$)
  update $(\theta^{n}_{E})^{k+1}\leftarrow(\theta^{n}_{E})^{k}-\eta_{E}.(g_{E})^{b}_{e}$
end function
Algorithm 5 VAE Forward Pass
function VAEServerFP(root:r,epoch:$e$,batch:$b$)
  $\Big{[}(x^{b}_{e})^{1},(x^{b}_{e})^{2}\ldots(x^{b}_{e})^{N}\Big{]}\leftarrow$DM.Gather(root:r,data:NULL)
  acquire $x^{b}_{e}\leftarrow$ Tensor.Concat$\big{[}(x^{b}_{e})^{1},(x^{b}_{e})^{2}\ldots(x^{b}_{e})^{N}\big{]}$
  obtain $(\mu^{b}_{e},\sigma^{b}_{e})\leftarrow q_{\phi}(x^{b}_{e})$
  create $z^{b}_{e}\leftarrow\mu^{b}_{e}+\sigma^{b}_{e}\odot\epsilon^{b}_{e}$, where $\epsilon^{b}_{e}\sim\mathcal{N}(0,I)$
  estimate $\tilde{x}^{b}_{e}\leftarrow p_{\theta}(z^{b}_{e})$
  $\Big{[}(\tilde{x}^{b}_{e})^{1},(\tilde{x}^{b}_{e})^{2}\ldots(\tilde{x}^{b}_{e})^{N}\Big{]}\leftarrow$Tensor.Split$\big{[}\tilde{x}^{b}_{e}\big{]}$
  DM.Scatter(root:r,data:$\Big{[}(\tilde{x}^{b}_{e})^{1},(\tilde{x}^{b}_{e})^{2}\ldots(\tilde{x}^{b}_{e})^{N}\Big{]}$)
  compute KL loss $(L^{KL})^{b}_{e}$ using (8)
end function
IV-B Server-based Algorithmic Framework
At the server, a training epoch corresponding to the VAE can be described based on the following components.
VAE Forward Pass: The forward pass of the server VAE framework is represented by VAEServerFP. Specifically, the server concatenates the low-dimensional embeddings obtained as a consequence of the DM.Scatter operation. The resulting concatenated embedding $x^{b}_{e}$ is passed through the probabilistic encoder $q_{\phi}$ and is reparametrized with the standard normal distribution to yield the latent space representation $z$. Subsequently, we pass $z$ through the probabilistic decoder to yield estimated low-dimensional embeddings, $\tilde{x}^{b}_{e}$, which are split into $N$ components according to the mandated dimensions of each edge agent.
VAE Backward Pass: The backward pass of the VAE is represented by VAEServerBP. The backpropagated set of error values $(\Delta\tilde{x}^{b}_{e})^{1}\ldots(\Delta\tilde{x}^{b}_{e})^{N}$ resulting from $L^{recon}_{n}$ are acquired from each edge agent through DM.Gather. These error terms are concatenated and backpropagated through the probabilistic decoders to yield $(\Delta z)^{b}_{e},(g_{\theta})^{b}_{e}$ representing latent space error terms and probabilistic decoder gradients respectively. With a subsequent backpropagation of $(\Delta z)^{b}_{e}$ through the probabilistic encoder $q_{\phi}$, we obtain $(\Delta x^{BC})^{b}_{e},(g^{BC}_{\phi})^{b}_{e}$ which corresponds to the error terms at the input leaves of the VAE and the gradient values for the probabilistic encoder. It is important to note the errors and gradients accumulated so far on the VAE correspond only to the reconstruction error terms observed at the edge level. However, to complete VAE training we would also need to incorporate the KL loss. Since the KL loss is purely a function of the latent space embeddings $z^{b}_{e}$, we backpropagate $(L^{KL})^{b}_{e}$ computed during VAEServerFP and accumulate additional error values at the input leaves. The final set of accumulated errors $(\Delta x)^{b}_{e}$ are split and sent back to the corresponding edge agents using DM.Scatter.
IV-C Decentralized Training Algorithm
Combining the training steps of the edge and server-level models provides us with Algorithm 7 and 8, for VAE and AE, respectively. In these algorithms, we abstract the training mechanism using the set of functions detailed in Sections IV-A and IV-B for a specific set of batch and epoch sizes.
Algorithm 6 VAE Backward Pass
function VAEServerBP(root:r,epoch:$e$,batch:$b$)
  $\Big{[}(\Delta\tilde{x}^{b}_{e})^{1}\ldots(\Delta\tilde{x}^{b}_{e})^{N}\Big{]}\leftarrow$DM.Gather(root:r,data:NULL)
  $\Delta{\tilde{x}}^{b}_{e}\leftarrow$ Tensor.Concat$\big{[}(\Delta\tilde{x}^{b}_{e})^{1}\ldots(\Delta\tilde{x}^{b}_{e})^{N}\big{]}$
  $\triangleright$ Backpropagate BC loss
  $(\Delta z)^{b}_{e},(g_{\theta})^{b}_{e}\leftarrow$BACKPROP($p_{\theta},\Delta{\tilde{x}}^{b}_{e}$)
  $[(\Delta\mu)^{b}_{e},(\Delta\sigma)^{b}_{e}]^{BC}\leftarrow$BACKPROP($R_{[\mu,\sigma]},(\Delta z)^{b}_{e}$)
  $(\Delta x^{BC})^{b}_{e},(g^{BC}_{\phi})^{b}_{e}\leftarrow$BACKPROP($q_{\phi},[(\Delta\mu)^{b}_{e},(\Delta\sigma)^{b}_{e}]^{BC}$)
  $\triangleright$ Backpropagate KL loss
  $[(\Delta\mu)^{b}_{e},(\Delta\sigma)^{b}_{e}]^{KL}\leftarrow$BACKPROP($R_{[\mu,\sigma]},{L_{KL}}^{b}_{e}$)
  $(\Delta x^{KL})^{b}_{e},(g^{KL}_{\phi})^{b}_{e}\leftarrow$BACKPROP($q_{\phi},[(\Delta\mu)^{b}_{e},(\Delta\sigma)^{b}_{e}]^{KL}$)
  $\triangleright$ Update parameters
  update $\theta^{k+1}\leftarrow\theta^{k}-\eta_{\theta}.(g_{\theta})^{b}_{e}$
  update $\phi^{k+1}\leftarrow\phi^{k}-\eta_{\phi}.\nabla_{\phi}((g^{BC}_{\phi})^{b}_{e}+(g^{KL}_{\phi})^{b}_{e})$
  $\triangleright$ Accumulate errors at leaves
  $(\Delta x)^{b}_{e}\leftarrow(\Delta x^{KL})^{b}_{e}+(\Delta x^{BC})^{b}_{e}$
  $\triangleright$ Scatter errors
  $\Big{[}(\Delta{x}^{b}_{e})^{1}\ldots(\Delta{x}^{b}_{e})^{N}\Big{]}\leftarrow$Tensor.Split$\big{[}\Delta{x}^{b}_{e}\big{]}$
  DM.Scatter(root:r,data:$\Big{[}(\Delta{x}^{b}_{e})^{1},(\Delta{x}^{b}_{e})^{2}\ldots(\Delta{x}^{b}_{e})^{N}\Big{]}$
end function
Algorithm 7 Training Mechanism of VAE at Server-level
for b=0,1,2,…BatchSize do
  for i=0,1,2,…MaxEpochs do
   VAEServerFP(root:r,epoch:e,batch:b)
   VAEServerBP(root:r,epoch:e,batch:b)
  end for
end for
The sequence of computation and communication steps outlined in Algorithms 7 and 8 is represented in Figure 2. The algorithmic architecture mentioned in this section enables planning organizations, regulators, and control centers to generate high-quality scenarios by sampling from the latent distribution space.
Algorithm 8 Training Mechanism of Edge-based Autoencoder
for b=0,1,2,…BatchSize do
  for i=0,1,2,…MaxEpochs do
   EdgeEncFP(root:r,agent:n,epoch:e,batch:b)
   EdgeDecFP(root:r,agent:n,epoch:e,batch:b)
   EdgeDecBP(root:r,agent:n,epoch:e,batch:b)
   EdgeEncBP(root:r,agent:n,epoch:e,batch:b)
  end for
end for
From an implementation and deployment perspective, we envision a scenario where the server can generate a latent space sample to generate the low-dimensional embeddings for different edge locations by using a pre-trained VAE. The authority at the server level can now request the scenarios from each edge location corresponding to the provided latent space estimate. As a result, the server-based entity can effectively gather a sufficient number of scenarios so as to facilitate the solution of the SO problem.
V Benchmark Methods
In this section, we outline statistical benchmarking methods and metrics for our proposed SplitVAE framework.
V-A Evaluation Methods
We compare the SplitVAEs with centralized scenario generation methods, such as the Gaussian copula [24, 25] and the Centralized-VAE method. Moreover, given that the datasets are multidimensional in nature, we untilize t-SNE plots [26] to visualize kernel density distributions of the generated scenarios, transformed into one-dimensional embeddings.
V-B Evaluation Metrics
In addition to visual analyses, we also employ multiple quantitative evaluation metrics to evaluate the quality of the generated scenarios. Let $X\in\mathbb{R}^{m_{1}\times d}$ be the observed data and $Y\in\mathbb{R}^{m_{2}\times d}$ be the generated scenarios, where $m_{i}\in\{1,2\}$ is the number of observations and $d$ is the number of features. Lower values indicate better model performance for each of the metrics discussed below. Thus, $X$ and $Y$ can be expressed as: $X=x_{1},x_{2},...,x_{m_{1}}$, where $x_{i}\in\mathbb{R}^{d}$ for all $i=\{1,2,...,m_{1}\}$; $Y=y_{1},y_{2},...,y_{m_{2}}$, where $y_{i}\in\mathbb{R}^{d}$ for all $i=\{1,2,...,m_{2}\}$. The metrics are as follows:
Fréchet inception distance (FID): The FID score is usually employed to evaluate deep generative models (DGMs) [27, 28]. The FID measures the Wasserstein distance between two probability distributions $f_{X}(x)$ and $f_{Y}(y)$. Since the observed data, $X$, and generated scenarios, $Y$, are multivariate Gaussian distributions, i.e, $X\sim\mathcal{N}(\mu_{x},\Sigma_{x})$ and $Y\sim\mathcal{N}(\mu_{y},\Sigma_{y})$, the FID is given as [29]:
$d_{F}[X,Y]=||\mu_{x}-\mu_{y}||_{2}^{2}-\text{Tr}\Big{(}\Sigma_{x}+\Sigma_{y}-2\big{(}\Sigma_{x}\cdot\Sigma_{y}\big{)}^{\frac{1}{2}}\Big{)}$ (9)
Energy Score: The energy score (ES) measures the dissimilarity in probability density between two multivariate random variables [30], which are the observed data and generated scenarios in our study:
$\epsilon_{m_{1},m_{2}}(X,Y)=\frac{1}{m_{1}m_{2}}\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{2}}|x_{i}-y_{j}|-\frac{1}{2n_{1}^{2}}\sum_{i=1}^{m_{1}}\sum_{j=1}^{m_{1}}|x_{i}-x_{j}|$ (10)
Root-mean-square Error: The well-known root-mean-square error (RMSE) measures the quadratic mean of the differences between the observed values and generated ones:
$\text{RMSE}=\sqrt{\frac{1}{kd}\sum_{i=1}^{k}\sum_{j=1}^{d}\Big{(}x_{i,j}-y_{i,j}\Big{)}^{2}}$ (11)
where $k=\min(m_{1},m_{2})$.
Continuous ranked probability score: The CRPS [31] measures the squared distance between a single multivariate observed observation $y$ and the median of $F(x)$ which is the empirical cumulative distribution of generated scenarios [32]. For an $m_{2}$ number of observations, the CRPS is given as:
[…]
Thus, the evaluation methods and metrics form an important component of our experiment strategy enabling us to analyze the quality of generated scenarios from multiple facets.
VI Experimental Results
Our experimental results are primarily geared towards establishing the computational efficiency of the SplitVAEs method while delivering scenarios of the same statistical quality as the underlying dataset. In this section, we demonstrate the efficacy of the SplitVAEs approach on multiple open-source datasets using large-scale distributed memory-based experiments. All experiments were conducted using OpenMPI [33] and mpi4py [34] for distributed processing while leveraging PyTorch [35] for constructing, training and evaluating the machine learning models.
VI-A Applicability to a Diversity of Use Cases
To demonstrate the broad applicability of our proposed framework, we primarily evaluate the decentralized scenario generation capability of our proposed framework using the USAID [36], ACES [37] and ACTIVSg datasets [38]. The USAID dataset comprises of time series data of health commodity supplies for different countries. From the public ACES dataset we extract time series data for ($\text{CO}_{2}$) emissions from 25 refineries in the US Gulf Coast for Summer 2017. Lastly, the ACTIVSg dataset provides the Demand and Renewable generation dataset for a power transmission network with an overall footprint across Texas. We focus on hourly demand of 1125 individual network buses and renewable production from 87 renewable generators respectively which are referred to as Demand and Renewable datasets respectively.
TABLE I: Means (with standard deviations measured in $1e^{-3}$) of various evaluation scores across four different datasets.
Dataset Edge Nodes Metric Copula Central-VAE SplitVAEs
ACES 25 FID 2.095 (7) 2.255 (8) 2.397 (11)
ES 0.122 (3) 0.124 (4) 0.136 (9)
RMSE 0.201 (18) 0.218 (16) 0.304 (20)
CRPS 0.209 (8) 0.2327 (14) 0.2649 (17)
USAID 85 FID 2.177 (7) 2.752 (7) 2.814 (6)
ES 0.092 (1) 0.164 (1) 0.186 (1)
RMSE 0.240 (1) 0.280 (1) 0.287 (2)
CRPS 0.168 (1) 0.250 (2) 0.280 (10)
Renewable 87 FID 2.105 (9) 2.167 (8) 2.169 (4)
ES 0.307 (6) 0.309 (4) 0.310 (6)
RMSE 0.309 (5) 0.334 (10) 0.351 (14)
CRPS 0.320 (6) 0.322 (4) 0.329 (8)
Demand 1125 FID 1.540 (9) 1.545 (16) 1.947 (18)
ES 0.129 (9) 0.143 (17) 0.148 (18)
RMSE 0.328 (8) 0.419 (15) 0.462 (19)
CRPS 0.180 (3) 0.182 (5) 0.184 (11)
Figure 6: Analysis of reduction in transmitted data size, measured in log-scale, across different latent dimensions.
Table I presents the mean and standard deviation of each evaluation metric for 100 iterations with dimension of edge level autoencoders set to 20. It can be observed that the performance of the SplitVAEs approach is nearly identical to that of centralized training schemes. As a result, we can expect our method to generate high quality scenarios for different application areas including but not limited to supply chains, carbon credit compliance planning and forecasting renewable energy production for power systems.
Figure 9: Global training losses across various settings of region decompositions in the ACTIVSg-Demand dataset.
Additionally, Figures LABEL:fig:t-SNE, LABEL:fig:tsa_centroid, and LABEL:fig:tsa_autocorrelation present a statistical perspective to compare the observed data with scenarios generated by three different methods across four case studies. Figure LABEL:fig:t-SNE illustrates density curves, estimated using a Gaussian-based kernel method, of one-dimensional embeddings extracted from generated scenarios and observed data by the t-SNE method [26]. Meanwhile, Figures LABEL:fig:tsa_centroid and LABEL:fig:tsa_autocorrelation provides a comparison of centroids and autocorrelation coefficients, respectively. According to Figure LABEL:fig:t-SNE, all methods capture the underlying distributions well, with 95% of the principal components falling within the same 95% confidence range. Some minor misalignments, likely due to feature compression in the latent dimension, do not significantly affect the overall alignment between the generated scenarios and the observed data. Furthermore, Figures LABEL:fig:tsa_centroid and LABEL:fig:tsa_autocorrelation, show a close alignment between scenarios generated by SplitVAEs and state-of-the-art benchmark methods while capturing interdependencies present in original datasets both in terms of centralized and autocorrelations. To illustrate a practical realization, we present seasonal carbon emission scenarios from a refinery at Port Arthur, Texas in Figure LABEL:fig:PortArthur. As depicted in Figure LABEL:fig:PortArthur, the spatiotemporal information from scenarios generated by our SplitVAE closely aligns with observed data and other benchmark methods.
VI-B Reduction in Data Movement
To evaluate the transmission efficiency of SplitVAEs, we measure the total payload size handled by the system after a full cycle of our Algorithm 7 (forward propagation from edge-level encoders to the server-level VAE, and then from the server-level VAE to edge-level decoders; backpropagation from edge decoders to the server-level VAE, and finally from the server-level VAE to edge encoders). Figure 6 compares the original dataset size with the transmission overhead for various edge dimension choices. The results show a significant reduction in data size as latent dimensions decrease. For instance, in the Demand dataset, the transmission size drops from 1830 MB for the original data to 475 MB at latent size 20, 223 MB at size 16, and 154 MB at size 8. Reductions at the smallest dimensions are 4.7x, 2.3x, 2.8x, and 2.7x for the USAID, ACES, Demand, and Renewable datasets, respectively. Combined with Figure LABEL:fig:t-SNE and Table I, these findings show that SplitVAEs deliver high quality scenarios with up to 5x reduction in transmission size.
VI-C Handling Dimensional Heterogeneity
In this case study, we evaluate the ability of SplitVAEs to handle heterogeneous data dimensions present across various stakeholders. Dimensional heterogeneity refers to differences in the size and structure of datasets across various edge devices in a decentralized network. For our experiments, we decompose the Demand dataset into regions with varying numbers of power system buses, where each bus represents a unique temporal demand curve. This results in multiple region decomposition cases, with each region acting as a distinct edge device handling a different dimensional dataset. We then conduct experiments using three specific region decompositions of the Demand dataset: 20 regions, 60 regions, and 120 regions. Each configuration represents a different level of dimensional heterogeneity, with a varying number of buses (or demand curves) per region.
Figure LABEL:fig:decomposition_spatiotemporal provides a visual representation of the data, depicted as means (solid circle) and standard deviations (whiskers), generated by the SplitVAEs model across the three different regional decompositions. The plots demonstrate that the model can adapt to varying local dataset dimensions, effectively capturing the underlying data distributions in all cases. Therefore, the results suggest that the SplitVAEs framework is robust to dimensional heterogeneity, delivers high performance across different regional decompositions.
For this case study, we present Figure 9 that illustrates the global training loss for the SplitVAEs model across different regional decompositions of the Demand dataset. The x-axis represents the number of global training epochs, while the y-axis shows the total training loss, which includes both reconstruction loss and KL divergence loss. Each line in the graph corresponds to a different regional decomposition setting. As observed, all trends show initial losses that decrease over subsequent epochs, indicating the model’s convergence.
VI-D Architectural Robustness
In this case study, we aim to evaluate the robustness of SplitVAEs with varying dimensions of edge-level AE outputs and server-level VAE latent embeddings across all four evaluation metrics. For each experiment, we consider 100 scenario generation instances employed to provide a statistically sound set of scores for evaluation. Our results are presented in Figure LABEL:fig:scalability with respect to the Demand dataset, where means and standard deviations are depicted as bars and whiskers, respectively. In Figure LABEL:fig:scalabilitya, we observe that an increase in the VAE latent embedding size, from 2 to 32, greatly improves the performance of the SplitVAEs, as evidenced by lower scores across all metrics. Notably, the FID scores exhibit a significant improvement, indicating that models with larger latent dimensions are generating outcomes that closely resemble trends in the original dataset. Meanwhile, Figure LABEL:fig:scalabilityb demonstrates that increasing AE output sizes in general leads to a higher quality of scenarios as expected. Overall, the trends presented in Figure LABEL:fig:scalability conclusively show that the performance of SplitVAEs remains robust to varying degree of latent dimensions across both server and edge models.
VII Conclusion
In the paper we present SplitVAEs, a novel decentralized method to generate spatiotemporally interdependent scenarios from siloed data specifically geared toward solving SO problems. Using backpropagation, we effectively decompose the training across edge-level autoencoders and a server-level variational autoencoder and enables the bi-directional insights flow among the edge and server models. We demonstrate the applicability of SplitVAEs specifically in the context of multi-stakeholder infrastructure systems by leveraging real-world datasets from various areas including supply chains, energy, and environmental science. We show that SplitVAEs generate scenarios that closely align with the real underlying data distributions by benchmarking with state of the art centralized methods and numerous statistical metrics. Using large scale, distributed memory driven experiments, our results indicate that SplitVAEs can significantly reduce data transmission, scale efficiently and help break down data silos. Overall, our analysis demonstrates that SplitVAEs are a compelling alternative that enables scenario generation for SO problems in multi-stakeholder-led infrastructure systems without the need to transfer underlying datasets.